OpenStudy (brenna):

Bob's speed was 20 mph faster than that of Lisa. Thus Bob could travel 240 miles in 2 hrs less than it took Lisa to travel 160 miles. Find the speeds and times of both.

OpenStudy (anonymous):

i wonder if this is right, because i get an ugly answer, so maybe i made a mistake

OpenStudy (anonymous):

yeah it is wrong

OpenStudy (anonymous):

bobs rate is x, lisa's is x - 20 bobs time is \[\frac{240}{x}\] lisa's is \[\frac{160}{x-20}\] and we know \[\frac{240}{x}=\frac{160}{x-20}+2\] that is better

OpenStudy (brenna):

i get the problem until i plug it into the equation Distance=Time x Rate. After i plug in the information it makes no sense.

OpenStudy (anonymous):

maybe it would make more sense if we tried it with numbers

OpenStudy (anonymous):

suppose lisa's speed is 40 then bob's speed has to be 60

OpenStudy (brenna):

i use bob's time x bob's rate = 240 miles then (lisa's rate -20) x (lisa's time +2) = 240 miles if i substitue the given information. after this point i get stuck

OpenStudy (bahrom7893):

lol satellite for some reason I read bobs as boobs 4 times.. and i was thinkin what the heck are these guys doing!

OpenStudy (anonymous):

at 60, bob would take 240/60= 4 hours lisa would take 160/40 = 4 hours. so their time would be equal

OpenStudy (anonymous):

lol

OpenStudy (brenna):

haha

OpenStudy (brenna):

how do you know bob's rate? it just says that it is 20 greater than lisa's

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

i don't know bob's rate, so i put it as x and so lisa's rate is x - 20 so far so good?

OpenStudy (anonymous):

then we are told information about time, and time is distance over rate so bob's time for 240 miles is \[\frac{240}{x}\] similarly lisa's time for going 160 miles is \[\frac{160}{x-20}\]

OpenStudy (brenna):

yea i got that and that bob's time is 2less than lisa's so lisa's rate is x+2

OpenStudy (anonymous):

it says bob goes 20 mph faster than lisa. so if bob goes 40 lisa goes 20, and if bob goes 60 lisa goes 40 and if bob goes x, lisa goes x - 20

OpenStudy (anonymous):

so now we know that bob's time is 2 hours less than lisa's time

OpenStudy (anonymous):

so \[\frac{240}{x}=\frac{160}{x-20}-2\]

OpenStudy (brenna):

yea

OpenStudy (anonymous):

now we have to solve for x somehow

OpenStudy (anonymous):

it is annoying that x is in the denomiator, but i do not think it can be helped

OpenStudy (mertsj):

 \[rt=160\] \[r=\frac{160}{t}\] \[(r+20)(t-2)=240\] \[rt-2r+20t-40=240\] \[160-2r+20t-40=240\] \[-2r+20t=120\] \[-2(\frac{160}{t})+20t=120\] \[\frac{-320}{t}+20t=120\] \[-320+20t^2=120t\] \[20t^2-120t-320=0\] \[t^2-6t-16=0\] \[(t-8)(t+2)=0\] \[t=8\]